OpenStudy (anonymous):

2. Terephthalic acid, an important raw material for making Dacron, a synthetic fibre, is made from para-xylene by the following reaction. C8H10 + 3 O2 ------------------------> C8H6O4 + 2 H2O para- special conditions terephthalic xylene acid How much terephthalic acid could be made from 154 grams of para-xylene in moles? In grams?

OpenStudy (chmvijay):

106 g of Para xylene produces 166 gram of terepthalic acid 154 gram of Para xylene produces =166*154 /106=241.6 gram if u want to convert in to moles = 241.6/166 = 1.45 moles